OpenStudy (anonymous):

When solving -1/8 (x + 35) = −7, what is the correct sequence of operations? A. Multiply each side by -1/8, add 35 to each side. B. Multiply each side by -1/8, subtract 35 from each side. C. Multiply each side by -8, subtract 35 from each side. D. Multiply each side by -8, add 35 to each side.

OpenStudy (anonymous):

\[-\frac{ 1 }{ 8 } (x + 35) = -7\]

OpenStudy (tylerd):

you wanna get rid of the -1/8 first, so what would you do?

OpenStudy (anonymous):

Multiply by -1/8, right?

OpenStudy (tylerd):

multiply by -8

OpenStudy (tylerd):

-8/-8 = 1

OpenStudy (tylerd):

1 times (x+35) = x+35

OpenStudy (anonymous):

So, multiply by -8 and subtract 35? Since you need to isolate the variable.

OpenStudy (tylerd):

yep

OpenStudy (anonymous):

Thank you so much!